### 30 Dec 2023 v2

# We assume the reader familiar with the definition of a partially ordered set, for which we will use the short form poset, and of the special case

1. Definitions and examples
We assume the reader familiar with the definition of a
partially ordered set, for which we will use the
short form poset, and of the special case
of a totally ordered set, also called a chain.
We will formally consider a poset to be an ordered
pair where is the underlying set
and the order relation; but when there no danger of
ambiguity, we shall write to mean
and write for
Posets will always be understood to be nonempty.
We will write etc., for
etc., and will use
etc. for the conjunctions of the
relations etc. with the relation
We shall use for the negation of
and likewise for the negation of etc..
Recall that a set-map between posets is called
isotone if
for all
and is called an embedding if it is isotone, and also satisfies
for all
Clearly, an embedding of posets is one-to-one; but a
one-to-one isotone map need not be an embedding.
A linearization of a partial ordering on a
set means a total ordering on which extends
in the sense that for if
then in other words, such that
the identity map of is an
isotone map
It is easy to verify that every partial ordering on a
set admits a linearization, and in fact that
Every partial ordering on a set is (as a set of
ordered pairs) the intersection of its linearizations.
(Idea of proof: Given such that show
that there is a strengthening of such that
Iterating this process infinitely many times if necessary, we get
a linearization of
Moreover, by our choice of the pair with that we start with,
we can insure that any prechosen order-relation that does not
hold under fails to hold in our linearization.
Hence looking at all linearizations, we have (1.3).)
By the product of two posets and one
understands the poset whose underlying set is
ordered so that
if and only if and
and analogously for products of larger families of posets.
(These are in fact products in the category
of posets and isotone maps, though we shall not use category-theoretic
language in this note.)
Thus (1.3) says
that every poset embeds in the product of the chains
obtained from by linearizing
Occasionally, I shall refer to a product of posets as their
‘direct product’, when this seems desirable for clarity.
We now come to the concept we will be studying in this note.
Definition 1.1.
The dimension of a poset is the least cardinal
such that
(i) the relation is an
intersection of total orderings on
equivalently, such that
(ii) is embeddable in a product of totally
ordered sets.
In the above definition, the implication (i)(ii) is clear.
To get (ii)(i) [14, Theorem 10.4.2])
first consider any isotone map
where is a totally ordered set.
For each regard the inverse image of in
as a subposet, and choose a linearization of
the partial ordering of that subposet.
We can now strengthen to a linear ordering
of by making if either
or and
Thus, given an embedding of in a product of
posets if we construct as above
from each of the projections a linearization
of we see that the intersection of
these linearizations will again be the partial ordering
We understand the product of the empty family of sets to
be a singleton.
Thus a poset has dimension if
and only if its underlying set is a singleton.
In the literature on dimensions of posets, condition (i)
above is generally
the preferred definition; but here we will more often use (ii).
This note will focus almost entirely on finite-dimensional
posets, though we will allow underlying sets
of these posets to be infinite.
In indexing finite families of maps etc.,
we will follow the set-theorists’ convention,
For a natural number,
For positive integers two important examples of posets of
dimension are:
The -cube i.e., the -fold
direct product where
denotes the poset with the ordering
and, for
The “standard example” whose underlying set
consists of elements
with the ordering that makes each less than every
other than and with no other order-relations among
these elements.
For can be identified with a subposet of by
identifying each with the -tuple that has
value in the -th coordinate and
elsewhere, and each with the -tuple that has
value in the -th coordinate and elsewhere.
is sometimes called the “crown” of dimension
This nicely fits the appearance of the diagram when
but not so nicely for higher
To see that (1.5) and (1.6) both have dimension
note first the inequalities
which are clear in view of the representation of as a subposet of
and of as a direct product of chains.
So it will suffice to show that the ordering on is
not an intersection of fewer than total orderings.
To see this, consider any family of total orderings of
whose intersection is the ordering of
Note that for each since in
that family of total orders must have at least one member
such that
I claim the orderings must be distinct.
Indeed, if for some
is the same as let us
call their common value
Interchanging the roles of and if necessary, we may
assume
Then hence under the intersection
of our family of orderings,
contradicting the definition of
So the orderings in our family are indeed distinct,
so so by (1.7), both (1.5)
and (1.6) indeed describe posets of dimension
For another class of examples of dimensions of posets, recall
that a poset in which every pair of distinct
elements is incomparable is called an antichain.
We claim that
Every antichain with more than one element has dimension
Indeed, if one chooses any total ordering of then the
intersection of that ordering and the opposite ordering is
the antichain ordering, so
and since has more than one
element, no single linear ordering makes it an antichain.
We remark that under an isotone bijection of posets
which is not an isomorphism,
the dimension may increase, decrease, or remain unchanged.
For instance, starting with an antichain of points,
a bijection onto the poset is an isotone map
that increases the dimension
from to while a linearization decreases the latter dimension
from to
On the other hand, a bijection from a -element antichain to
the poset is isotone and leaves the dimension, unchanged.
Let us note one more elementary poset construction
(a generalization of one we applied to chains
in proving the equivalence of the two conditions
in Definition 1.1).
Given a poset and for each a poset
we may define the set
and partially order this lexicographically; that is, by defining
if and only if either
or and
It is not hard to verify that for this construction,
is the supremum of
In the verification, one takes a representation
for as in Definition 1.1(i), and
representations of the as in Definition 1.1(ii)
(which may, but do not need to satisfy the stronger condition (i)).
By repeating mappings if necessary, one can assume that the number of
chains used in each of these representations is the supremum
indicated above.
One can then easily combine these to get an embedding
of in a product of that number of chains.
2. Our main result, and some consequences
To lead up to our main result, let us first note
that for any two posets and (which, we recall,
are required to be nonempty), we must have
The properties of vector-space dimension
suggest that the second inequality above
should be equality; but the two concepts of dimension
are not alike in that respect.
To see an interesting
way that equality can fail, consider a poset of
dimension represented as a subposet of a
product of chains,
and consider a nontrivial chain
Suppose that for each we let ordered
lexicographically – intuitively, the chain gotten from by
replacing each element by a miniature copy of
It is easy to see that the map
taking to
is one-to-one and isotone.
Can it fail to be an embedding?
In general, yes.
If in and in then in the
product poset
the elements and are by definition incomparable;
but if it happens that for all we have
then under the ordering described above,
However, if the subposet
has the property that whenever two elements satisfy
the elements
agree in at least one coordinate,
then looking at that coordinate, we see that we do get
incomparability between the indicated elements of
so is an embedding, so
we indeed have the desired result,
The condition that every pair of comparable elements
of agree
in at least one coordinate may seem unnatural, but
for it is easy to see that it holds in the poset
regarded as a subposet of (sentence following (1.6)).
In fact, more is true there: Every such pair of elements
agree in one coordinate where their common value is
(the -th coordinate),
and one where their common value is (the -th);
and the above construction can be adapted to show, as a
consequence, that taking the product of with two
chains does not increase its dimension.
The next result gives this argument in detail for a
still more general situation, that applies to products
of with possibly more than two chains.
Note, however, that in that theorem, is not assumed to
be the dimension of (as it is in the
above example), but simply an integer such that
is a subposet of a product of chains with certain properties.
Theorem 2.1.
Let and be nonnegative integers, a
-tuple of chains, which for notational convenience we will
assume pairwise disjoint, a subposet of
and an -tuple of pairwise disjoint
subsets of such that
For every comparable pair of elements
in and every there exists such
that the -th coordinates of and are equal,
and their common value is a member of
Then for any -tuple of chains we have
Proof.
For notational simplicity, we may assume all the are equal
to a common chain and that
Indeed, given structures as in the statement of the theorem,
we may embed all the in a common chain and if we
prove (2.3) with the all replaced by
this will imply the same inequality for the original
choices of
Likewise, if we enlarge some of the
(still keeping them disjoint) so that their
union becomes the whole set
then the hypothesis
about comparable pairs of elements that was assumed
for the original choices of remains true for these
enlarged sets.
To obtain (2.3), we need an embedding of the poset
on the left-hand side, which we can now write
in a product of chains
For this purpose we define
the set-theoretic product
of and ordered lexicographically, i.e., so that
if and only if
or and
Also, since the partition
we can set the notation
For each will denote the
unique value such that
We now define a map as follows.
For
and let
be the element of whose
-th coordinate is for each
We wish to show that is an embedding of posets.
First, is isotone, i.e.,
If in then
in
This follows easily from (2.5) and (2.7).
To complete the proof that is an embedding, we need to show that
If then
This breaks down into two cases.
Suppose first that
In that case, for some we have
i.e., and
looking at the -th coordinates of and
namely and
we see from the lexicographic ordering of
that the former is the latter,
giving the conclusion of (2.9).
If, on the other hand,
but
then in view of the second inequality above,
we may choose a such that
Now by the first inequality of (2.11) and
the hypothesis (2.2), there is some such that
Note that
and have -th terms
and respectively.
Since the -coordinates of the above two -th terms
are the same by (2.13), the order-relation
between them is
that of the -coordinates, which satisfy (2.12).
This completes the proof of (2.9), as required.
Remark:
I came up with the above result after pondering [4],
which showed by an explicit construction that
i.e., is equal to
The next corollary includes that case.
Corollary 2.2.
Suppose and
is a family of chains, each
having a least element and a greatest element
and is a subposet of consisting of
elements each of which has at least one coordinate of
the form and at least one of the form
Then for any two chains and we have
In particular, if
and is any subposet of
containing the subposet (e.g., if
then for any two chains and
Proof.
In the context of the first assertion,
given since has
at least one coordinate of the form
must also
have as its -th coordinate; and similarly,
since has a coordinate must agree with
in that coordinate.
Applying Theorem 2.1
with
we get (2.15).
The second statement follows because when each is
the -element set the exclusion
of and from forces every element
to have at least one coordinate of
the form and one of the form
and the assumption that contains which
we saw following (1.6) has dimension turns the
inequality (2.15) into equality.
If, instead of assuming that every element
has at least one -coordinate and
at least one -coordinate, we only assume one of these
conditions, it is easy to check that the analogous reasoning
gives a conclusion half as strong:
Corollary 2.3.
Suppose and is a family of chains each
having a least element,
and is a subposet of consisting of
elements each of which has at least one coordinate of the form
Then for any chain we have
In particular, if is any subposet of
containing the subposet (e.g., if
then for any chain we have
The analogous statements hold with least elements and
everywhere replaced by greatest elements and ∎
To get examples of Theorem 2.1
with we shall use a different way of choosing subsets
of based on the subscript rather
than the distinction between greatest and least elements of
We will need a bit more notation.
For every positive integer and every pair of integers
with
Let denote the subset of consisting
of those elements in which the number of coordinates
of the form is either or
(Thus, for
Corollary 2.4.
For every positive integer every integer with
and every family of chains with
we have
Proof.
Let us define subsets of
by
Condition (2.20) guarantees that all of the
are contained in
Now if are elements of then since the
cardinalities of the subsets of on which they assume values of
the form differ by at most they must agree
in all but at most one coordinate.
Since each contains both and for
two values of one of those values of must have
the property that and agree on the -th coordinate.
Hence these subsets satisfy the hypotheses
of Theorem 2.1, completing the proof.
If the posets had, like
dimension
then the above result would show that for every there exist
finite posets that don’t change their dimension on
taking a direct product with chains.
However, such subposets of generally have dimension less
than
There are many results in the literature obtaining
bounds on the dimensions of subposets of
cf. [5], [6], [7, Theorems III and V],
[8, Theorem 7.1], [10],
[11, Theorems 7, 10, 12, 13], and [16, Theorem 2].
Let us obtain one such result here, as another consequence
of Theorem 2.1.
(In the statement and proof, we shall not need to look
at the disjoint union of the factors of
hence we shall not, as in the preceding
corollaries, treat these as disjoint sets
but as the same set
Corollary 2.5.
Let be a positive integer, define
and let
(where by we mean the -tuple whose
first coordinates are and whose last
coordinates are and by the
same with the roles of and reversed).
Then
Hence the same upper bound holds for the dimension of
any subposet of
in particular, for any poset of the form
(see (2.19))
with
Proof.
I claim that as a subset of
in other words, that the first coordinates
of every element of contain at least one and
at least one
To see the former condition, note that by (2.23),
itself must have at least two coordinates
If none of these were among the first coordinates,
this would force but that element is
excluded in (2.24).
The second condition follows in the same way.
Now by Corollary 2.2, with in place of
and each taken to be we see that the direct
product on the right-hand side of (2.26) has
dimension hence the same is true of its
subposet proving (2.25)
The final sentence of the corollary follows immediately.
3. An old result, and questions old and new
Having seen that dimension of posets is not in
general additive on direct products, it is striking that in an important
class of cases, it is.
The theorem below was proved, for products over index sets of arbitrary
cardinality, in K. Baker’s unpublished undergraduate
thesis [1].
That proof is summarized in [8]
for pairwise products (from which the case of arbitrary
finite products follows), and assuming the dimensions finite.
I give below a version of the proof (also for pairwise products of
finite-dimensional posets) which I find easier to follow.
Theorem 3.1.
([1, p.9, Property 3], [8, p.179, last 11 lines])
Let and be finite-dimensional posets, each
having a least element
and a greatest element
Then
Proof.
Clearly holds in (3.1), so it suffices
to show
For this we shall use version (i) of Definition 1.1,
and show that if we have an expression for the partial order
of as an intersection of
total orderings, on
then we can split that family into
two disjoint subsets, such that a certain natural map of into
the product of chains determined by one of those subsets is
an embedding, as is a map of
into the product determined by the other.
Thus the former subset must consist of
orderings, and the latter of orderings, so
as required.
The trick to finding this partition is to look at the
relative order, under each of the orderings (3.2),
of the elements and of
Reindexing those orderings if necessary, we may assume that
for some
for while
for
Let us now show that the map
given by
is an embedding.
Since (3.4) is clearly isotone,
it remains to show that if satisfy
then the corresponding condition holds on
the images of and under (3.4).
To get this, note that under the product ordering on
the relation (3.5) implies that
Hence we can find some such that
Now for all note that
contradicting (3.6);
so an satisfying (3.6) must be
Since (3.6) implies
this completes the proof that (3.4) respects
hence is an embedding.
The analogous argument shows that the map
given by gives an embedding of
As noted in the first paragraph, this proves (3.1)
Though we have seen that (3.1)
does not hold without the assumptions
that and have upper and lower bounds, the deviations
from equality in all known examples are
leading to the longstanding open question:
Question 3.2.
(i) If and are finite-dimensional posets,
must ?
In particular,
(ii) If is a finite-dimensional poset, and
a positive integer such that
must ?
In [8, last line of p.179 and top line of p.180]
it was conjectured that can exceed
by at most the number of members
of the set that do not have both a
greatest and least element.
Thus, the case where that number is is
Theorem 3.1, while the case where there is no restriction
on or corresponds to Question 3.2(i) above.
The case of that conjecture
where that number is however, turned out to
be false: from Corollary 2.2 we see that
for any
so exceeds
by
though has both greatest and least elements.
However, one might ask about different intermediate cases
between those of Theorem 3.1 and
Question 3.2(i),
suggested by Corollary 2.3, and noted in
Question 3.3 below.
C. Lin [11] poses part (iii) of that
question, though Theorems 10 and 11 of that paper
suggest the stronger implications of parts (i) and (ii).
Question 3.3.
(i) If and are finite-dimensional posets
such that has a least element and
has a greatest element,
must ?
(ii) If and are finite-dimensional posets
each of which has a least element,
or each of which has a greatest element,
must ?
And finally, a possible implication weaker than either of the above two:
(iii) [11, p.80] If and are
finite-dimensional posets such that has a least or a
greatest element, and has both,
must ?
Turning back to Question 3.2, I wondered whether
the result of Theorem 3.1, where no dimensionality
was lost in forming a product of posets, could be related to the
hypothesis that the sets of minimal and of maximal elements of each
factor were singletons, and hence had dimension
If so, could one use the fact that for arbitrary finite
and the sets of minimal and maximal
elements are antichains, hence have dimension
to get a positive answer to Question 3.2 for
finite posets?
But I see no way to adapt the proof of Theorem 3.1.
Moving on to other questions, recall that
in Theorem 2.1, the result did not depend on
the lengths of the chains
This suggests
Question 3.4.
(Cf. [13], [15, Conjecture 1])
If is a finite-dimensional poset, and any chain of
more than one element, must ?
The next question at first seemed “obviously” to have
an affirmative answer; but I see no argument proving it.
Question 3.5.
If is a finite-dimensional poset,
and
must ?
Much more generally (but much more vaguely), we may ask
Question 3.6.
Given finite-dimensional posets
if we know their dimensions,
and those of and
what can we say about ?
(Anything more than that it is greater than
or equal to the dimensions of each of
the pairwise products, and less than or equal to the values of
for ?)
One example of “misbehavior” on this front:
Suppose is an antichain of more than one element, and
Now and are all
antichains, hence all have dimension
but and
have the forms and
and from (1.11) we see that they
have dimensions and respectively.
So the dimensions of three
posets, and of their pairwise direct products, do not
determine the dimension of the product of all three.
4. Absorbency
The following concept might be helpful
in studying questions of the sort we have been considering.
Definition 4.1.
If is a finite-dimensional poset, let us define its
absorbency to be
the largest natural number such that
for every
-tuple of chains
As an example,
For all
Here follows from the case of (2.16).
To get the reverse inequality, we will need to call on a result
from the literature.
Note that if were then since
the value of would be
However, Trotter [16, Theorem 2]
shows that for all
An immediate property of absorbency is
since for any the product of with, say, will
contain a copy of and
hence have dimension greater than that of
Here are some other easy properties.
Lemma 4.2.
Let and be finite-dimensional posets.
If then
and
The case where neither the hypothesis of (4.4)
nor the corresponding inequality with the roles of
and reversed holds is covered by
If
then
Proof.
In (4.4), the first conclusion holds
because embeds in a product of chains,
which can “absorb” without increasing its dimension.
To get the final inequality of (4.4),
we must show that the dimension
of is not increased on multiplying it by a
product of chains.
Let us write such a product as where
is a product of chains,
and a product of chains.
Then if we write as
we see that
has dimension (by our choice of hence
has (by our choice of dimension at most
so its
product with has dimension which
we have noted equals so we have
indeed shown that the dimension
of has not been increased, as required.
In the conclusion of (4.5), the left-hand inequality
is immediate.
The right-hand inequality is equivalent to saying that
is bounded above by minus either of
and so by symmetry it suffices to
prove the bound involving
If we embed in a product of nontrivial chains,
and break this into a product of chains
and a product of chains, then
we have
as desired.
Question 4.3.
Under the hypothesis of (4.4),
does equality always hold in the final inequality of that implication?
If we ask the same question about
the final inequality of (4.5), the peculiarities
noted in the paragraph following Question 3.6 make trouble.
But those involved posets constructed using
antichains, and I know of no cases not arising in that way.
To express this precisely, let us set up some terminology.
Definition 4.4.
We shall call a poset disconnected if it satisfies
the following equivalent conditions.
(i) can be written as a union of subposets
and such that all elements of
are incomparable with all elements of
(ii) can be written as a union
of a family of more than one subposets
such that for all elements of
are incomparable with all elements of
(iii) is isomorphic to a poset
constructed as in (1.9) and (1.10)
with the index-poset (there called an antichain
of cardinality
On the other hand,
a poset will be called connected if it
is not disconnected; equivalently, if the equivalence
relation on generated by the relation
is the indiscrete equivalence relation.
Question 4.5.
If the posets and in (4.5)
are connected, must equality
hold in the final inequality of the conclusion thereof?
Since (1.11) gives us a way of computing the
dimension of a disconnected poset from the dimensions of its
connected components, a positive answer to
Question 4.5 (together with the
first assertion of (4.4)) would
allow us to compute exactly the dimension of a product of
two arbitrary finite-dimensional posets, given the dimensions and
absorbencies of their connected components.
It would, in particular, imply positive answers to
Questions 3.4 and 3.5.
But as long as we do
not know a positive answer to Question 4.5,
here are some further points worth considering:
If Question 3.4 has a negative answer, one could
consider variants of the absorbency concept, depending on
the lengths of the chains involved; though this is likely
to get messy.
Assuming Question 3.4 has a positive answer
but Question 3.5 does not,
one might define the “eventual absorbency”
of a finite-dimensional poset as the supremum, as
of
Given a poset and a nonnegative integer
about which we know that one might define
the absorbency of relative to to
be the largest such that the product
of with every -tuple of chains has dimension
So, for instance, given finite-dimensional posets and
the absorbency of relative to
will be at least the sum of the absorbencies of and of
5. Another function related to dimension, and a few from the literature
Let us call a poset bounded if it has a greatest element
and a least element.
(I wish I could think of a better term for
this condition, since it has very different properties from
the familiar sense of “bounded”.
E.g., a subposet of a bounded poset is not, in general, bounded.
I would welcome suggestions!
However, “bounded” is used in this way in places in the literature;
e.g., [8, p.179, 12th line from bottom].)
This allows us to define
another function which is useful in studying dimensions.
Definition 5.1.
For a finite-dimensional poset, the
bounded dimension of
denoted will be defined to be
(i) the greatest integer such that contains a
bounded subposet satisfying
equivalently,
(ii) the maximum, over all pairs of elements
in of
Since every bounded subset of is contained in an
interval we see that
the two versions of the above definition are equivalent.
Because of Theorem 3.1,
this function behaves very nicely under direct products:
Lemma 5.2.
For finite-dimensional posets and
Proof.
For elements of let us write
regarded as
a subposet of
Thus, by version (ii) of the definition of bounded dimension,
is the greatest of the
values for
in
But by the order-structure of a direct product,
is isomorphic to
and by Theorem 3.1 the dimension of this
product is
Taking the maximum over all pairs and
we get
Clearly, for any poset
However, the two sides of (5.3) can be far from equal, as
can be seen by taking for large
Then the left-hand side of (5.3)
is easily seen to be while the right-hand side is
The concept of bounded
dimension allows us to get an upper bound on absorbency:
Indeed, if we take the product of with more
nontrivial chains than the number on the right,
then since contains a bounded set of dimension
and each of the chains contains
a nontrivial bounded chain, the whole product will have dimension
Thus, to be “absorbed”, a family of
nontrivial chains can have at most the number of members
shown on the right in (5.4).
The two sides of (5.4) can also be far from equal.
For instance, for where
we have seen in (4.2) that
the left-hand side of (5.4) is
On the other hand, the right-hand side is
From (5.4) we get a strengthening
of (4.3); namely, we can add
to that inequality the observation,
The only finite-dimensional posets for which
are the antichains,
since a poset that is not an antichain contains a copy of
a bounded poset of dimension
Let us note some examples regarding
the inequalities in the conclusion of (4.5) if we
do not make the connectedness assumption of Question 4.5.
(These translate the “misbehavior” noted in the last
paragraph of §3.)
Let be a -element antichain, and
as usual, the -element chain.
Then has dimension and absorbency
while the poset has dimension and, by (5.4),
absorbency
Hence by (4.4), has dimension
and absorbency at least so by (5.4) its
absorbency must be exactly
The analogous considerations show that
i.e., again has
dimension but has absorbency
Hence taking
the inequalities in the conclusion of (4.5)
become so the first is equality
and the second a strict inequality.
On the other hand, if we take
then the corresponding considerations show that
this poset again has dimension but has absorbency while
has dimension
so the inequalities become
with the first strict and the second an equality.
(If, for a further comparison, we take
we find that the inequalities in question are
hence are both equalities.)
Trotter, in [16, Conjecture 2] suggested that
for all there exist posets of dimension
such that also has dimension
Reuter [15, Theorem 13] showed that no such exists for
but it is conceivable that there exist such for higher
Such an example would imply a negative answer to
Question 4.5, in view of (5.4).
Returning to the bounded-dimension
function, it might be of interest to look at the variants
of in which bounded is replaced by
bounded above, respectively, bounded below.
I will end this section by sketching some
other functions related to dimension that appear in the literature.
A striking example, studied in [2], [12],
[18] is based on the concept of a Boolean
representation of a poset
Here one considers families of total orderings
of which are not assumed to be strengthenings of the
ordering of but merely to have the
property that the order-relation or incomparability between
elements in can be expressed in terms
of which of the given orderings have and which have
For example, the order relation of
can be so described in terms of the four total orderings
Namely, in if and only
if on the one hand, precedes
in (5.6) but follows it in (5.7) (which together
guarantee that has the form and the form
and, further, the relative order of and
is the same in (5.8) as in (5.9) (which guarantees
that they are not a pair of the form
This is called a Boolean representation of in
terms of the four orderings (5.6)-(5.9).
The Boolean dimension of a poset
is the least such
that has a Boolean representation in terms of total orderings.
This number is always but often strictly less;
e.g., the above example shows that for all
In [18] it is shown that agrees with
when the latter is
(In the formal definition of a Boolean representation
of a poset in terms of total orderings
of one begins by mapping
to by sending to the -tuple which
has in the -th position if and only if
The ordering of is then
determined by a function
such that if and only if
takes the -tuple determined by to
For instance, in the above example describing is
the function on taking the value only
at the two -tuples and
Here the condition that for a -tuple to be taken to
its first entry must be and its second entry
translate the conditions that
must precede in (5.6) but not in (5.7);
and the condition that the last two entries of the -tuple
be equal translates the condition that and must occur
in the same order in (5.8) and in (5.9).
The term “Boolean dimension” refers to the fact that a map
can be expressed by a Boolean word
in variables.
Incidentally,
I think that in the definition of a Boolean representation, it would be
desirable to allow in which case the Boolean dimension of an
antichain would be determined by the function
sending the unique element of to meaning
that no pairs satisfy
Another variant of dimension described in [2] is
the local dimension, the least such that
there exists a family of linearizations of subposets of
such that each member of appears in of these linearized
subposets, and such that for every pair
of distinct elements of there are
enough linearized subposets containing both so that the
relationship between and in (an order-relation, or
incomparability) is the relation between their images in the
product of these linearizations.
As with the Boolean dimension, this is and
considerably less for the (in this case always
Still another variant of
dimension: If instead of looking at lists of linearizations
of such that every relation in is realized
in at least one member of our list, and letting the dimension of
be the least cardinality of such a list, one can look at
lists of linearizations with a “weight”,
a positive real number, attached to each, such that
in if and only if the sum of the weights of the
linearizations for which is at least
and define the fractional dimension of to be the infimum,
over all weighted lists which determine the ordering of
in this way, of the sum of the weights of the listed linearizations.
(The term “weight” is not used in the literature; it is my way
of giving an intuitive description of the definition.)
See [9, p.5, and references given there].
([9] also introduces several dimension-like
functions specific to the type of posets there named
“convex geometries”.)
Less exotic invariants of posets considered in [2] and
elsewhere are
the height, i.e., the supremum of the cardinalities of chains
contained in and the width, the supremum of the
cardinalities of antichains in
The function associating to a poset
the largest such that contains a subposet isomorphic
to is studied, under the name
in [3], [11, §3], [17, §5.2.1].
The function though usually (as in this note) simply called
the dimension of is sometimes, as in the title of [2],
called the Dushnik-Miller dimension, to distinguish
it amid the plethora of variant concepts.